OpenStudy (anonymous):

What would be the solution to -3(x+5) = 3(x-1)

OpenStudy (bahrom7893):

-3x-15 = 3x-3 -3x-3x = -3 + 15 -6x = 12 x = -2

OpenStudy (anonymous):

thanks